PLEASE explain how to get to the answer too.

Retail Inventory Method

Harmes Company is a clothing store that uses the retail inventory method. The following information relates to its operations during the year:

Cost Retail
Inventory, January 1 $29,200 $41,500
Purchases 66,200 101,700
Markups (net) 2,200
Markdowns (net) 200
Sales 86,500

Required:

1. Compute the ending inventory by the retail inventory method for the following cost flow assumption: FIFO. Round the cost-to-retail ratio to three decimal places. If necessary, round dollar amounts to the nearest whole dollar.

HARMES COMPANY
Calculation of ending inventory by retail inventory method
FIFO
Cost Retail
Purchases $ $
Markups (net)
Markdowns (net)
$ $
Add: Beginning inventory
Goods available for sale $ $
Less: Sales
Ending inventory at retail $
Ending inventory at cost $

2. Compute the ending inventory by the retail inventory method for the following cost flow assumption: Average cost. Round the cost-to-retail ratio to three decimal places. If necessary, round dollar amounts to the nearest whole dollar.

HARMES COMPANY
Calculation of ending inventory by retail inventory method
Average Cost
Cost Retail
Beginning inventory $ $
Purchases
Markups (net)
Markdowns (net)
Goods available for sale $ $
Less: Sales
Ending inventory at retail $
Ending inventory at cost $

3. Compute the ending inventory by the retail inventory method for the following cost flow assumption: LIFO. Round the cost-to-retail ratio to three decimal places. If necessary, round dollar amounts to the nearest whole dollar.

HARMES COMPANY
Calculation of ending inventory by retail inventory method
LIFO
Cost Retail
Beginning inventory $ $
Purchases
Markups (net)
Markdowns (net)
$
Goods available for sale $ $
Less: Sales
Ending inventory at retail $
Ending inventory at LIFO cost
Beginning layer (as stated in data) $
New layer
Total $

4. Compute the ending inventory by the retail inventory method for the following cost flow assumption: Lower of cost or market (based on average cost). Round the cost-to-retail ratio to three decimal places. If necessary, round dollar amounts to the nearest whole dollar.

HARMES COMPANY
Calculation of ending inventory by retail inventory method
Lower of Cost or Market (based on average cost)
Cost Retail
$ $
$ $
$
$
Answer #1

Answer -

1. FIFO

Cost Retail
Purchases $66200 $101700
Markups (net) $2200
Markdowns (net) ($200)
$66200 $103700
Add: Beginning inventory $29200 $41500
Goods available for sale $95400 $145200
Less: Sales ($86500)
Ending inventory at retail $58700
Ending inventory at cost $37473

Cost to retail ratio = ($66200 / $103700) = 0.638

So, Ending inventory at cost = $58700 * 0.638 = $37473

2. Average cost

Cost Retail
Beginning inventory $29200 $41500
Purchases $66200 $101700
Markups (net) $2200
Markdowns (net) ($200)
Goods available for sale $95400 $145200
Less: Sales ($86500)
Ending inventory at retail $58700
Ending inventory at cost $38567

Cost to retail ratio = ($95400 / $145200) = 0.657

So, Ending inventory at cost = $58700 * 0.657 = $38567

3. LIFO

Cost Retail
Beginning inventory $29200 $41500
Purchases $66200 $101700
Markups (net) $2200
Markdowns (net) ($200)
$66200 $103700
Goods available for sale $95400 $145200
Less: Sales ($86500)
Ending inventory at retail $58700
Ending inventory at cost
Beginning layer (as stated in data) $29200
New layer $10980
Total $40180

Cost to retail ratio (Beginning layer) = ($29200 / $41500) = 0.704

So, Ending inventory at cost (Beginning layer) = $41500 * 0.704 = $29200

And

Cost to retail ratio (new layer) = ($66200 / $103700) = 0.638

So, Ending inventory at cost (new layer) = $17200 * 0.638 = $10980

4. Lower of cost or market (based on average cost)

Cost Retail
Beginning inventory $29200 $41500
Purchases $66200 $101700
Markups (net) $2200
$95400 $145400
Markdowns (net) ($200)
Goods available for sale $145200
Less: Sales ($86500)
Ending inventory at retail $58700
Ending inventory at cost

Cost to retail ratio = ($95400 / $145400) = 0.656

So, Ending inventory at cost = $58700 * 0.656 = $38514